Solve the following triangles. Identify the case # in each given
triangle.

6.) Given parts:

a= 173 b= 267 c=412
1
Expert's answer
2020-11-10T18:55:41-0500

Here is the sketch of the triangle;


We are given all the three sides. We need to calculate the three angles A, B, and C.

Here we apply cosine rule.

"a^2 = b^2+c^2 - 2bccosA"

And this can be re-written as;

"cosA = \frac{b^2 + c^2 - a^2}{2bc}"


"cos A = \frac{267^2+ 412^2 - 173^2}{2 × 267 × 412}"


"=\frac{71289+169744-29929}{220008}"


"=\frac{211104}{220008}"


"= 0.95952874441"


"\therefore cosA =0.95952874441"

"A = cos^{-1} (0.95952874441)"

"= 16.36 \u00b0"


To find angle B, we again apply cosine rule

"b^2= a^2 +c^2 - 2ac cos B"

Which can also be re-written;

"cos B= \frac{a^2+c^2-b^2}{2ac}"


"cos B= \frac{173^2+412^2-267^2}{2 × 173 × 412}"


"=\frac{29929+169744-71289}{142552}"


"=\frac{128384}{142552}"


"=0.90061170661"


"B = cos^{-1}(0.90061170661)"


"B= 25.76 \u00b0"


To find the final angle C, we apply the property: 'The sum of the three interior angles of a triangle add up to 180°'


"\therefore C = 180 ° - (A+B)"


"= 180 \u00b0 - (16.36 \u00b0 + 25.76 \u00b0)"

"=180 \u00b0 - 42.12 \u00b0"

"=137.88 \u00b0"


Therefore, the three angles are;

"A= 16.36 \u00b0"

"B = 25.76 \u00b0"

"C=137.88 \u00b0"